# Demystifying Spectral Feature Learning for Instrumental Variable Regression

### v2

@@ Abstract
We address the problem of causal effect estimation in the presence of hidden confounders, using nonparametric instrumental variable (IV) regression. A leading strategy employs spectral features - that is, learned features spanning the top eigensubspaces of the operator linking treatments to instruments. We derive a generalization error bound for a two-stage least squares estimator based on spectral features, and gain insights into the method’s performance and failure modes. We show that performance depends on two key factors, leading to a clear taxonomy of outcomes. In a good scenario, the approach is optimal. This occurs with strong spectral alignment, meaning the structural function is well-represented by the top eigenfunctions of the conditional operator, coupled with this operator’s slow eigenvalue decay, indicating a strong instrument. Performance degrades in a bad scenario: spectral alignment remains strong, but rapid eigenvalue decay (indicating a weaker instrument) demands significantly more samples for effective feature learning. Finally, in the ugly scenario, weak spectral alignment causes the method to fail, regardless of the eigenvalues’ characteristics. Our synthetic experiments empirically validate this taxonomy.
@@ 1 Introduction
$\bullet$ We empirically validate the “good–bad–ugly” taxonomy through synthetic experiments designed to isolate the effects of spectral alignment and singular value decay on generalization error.
@@
Features can be learned in various ways. One line of work learns them jointly with the regression objective, as in [15, 22, 31, 37], who propose end-to-end frameworks that simultaneously optimize over the feature representations and second-stage parameters. In contrast, [34, 36] propose to learn features that reflect the eigenstructure of $\mathcal{T}$, independently of the outcome $Y$, using a spectral contrastive loss. This may appear suboptimal, since incorporating $Y$ could, in principle, yield more informative representations, however it avoids the optimization difficulties inherent to nonconvex, joint objectives involving all three variables $(X,Y,Z)$. This decoupled spectral approach was shown in certain experimental settings to outperform end-to-end alternatives, as demonstrated in [34]. Our work places this empirical finding in context, since it likely arises due to the experiments in that work satisfying the “good” regime in the present work. We provide further discussion in Appendix B.
Structure of the paper. Section 2 introduces the notation and preliminaries. Section 3 reviews the Sieve 2SLS estimator and presents generalization bounds highlighting the role of spectral features. Section 4 shows how to learn such features via a contrastive loss and connects this to the method of [34]. Section 5 provides synthetic experiments that validate our theoretical taxonomy. Proofs are deferred to the Appendix.
@@ 2.2 On the Conditional Expectation Operator: NPIV as an Ill-posed Inverse Problem
The existence part is essential: without it, there is no structural function $h_{0}$ satisfying the model in Eq.˜1. While injectivity can in principle be relaxed, this comes at the cost of recovering a different object. Specifically, if $h_{0}$ is not identifiable, one may instead recover the minimum-norm solution to Eq.˜2, namely $h_{*}\doteq\mathcal{T}^{\dagger}r_{0}$ [14]. [^2] This corresponds to the unique element in the pre-image of $r_{0}$ under $\mathcal{T}$ that minimizes the $L_{2}$-norm and is equivalently characterized as the unique element of the set $\mathcal{T}^{-1}(\{r_{0}\})\cap\mathcal{N}(\mathcal{T})^{\perp}$ [2]. We assume injectivity henceforth to simplify the discussion.
@@
Proposition 1.
Under Assumption˜2, $\mathcal{T}$ is a Hilbert–Schmidt operator and thus compact.
This result is well known in the inverse problems literature; we provide a proof in Appendix A.1.
@@
where $\{u_{i}\}\subset L_{2}(Z)$ and $\{v_{i}\}\subset L_{2}(X)$ are orthonormal systems (left and right singular functions, respectively) spanning $\mathcal{N}(\mathcal{T})^{\perp}$ and $\overline{\mathcal{R}(\mathcal{T})}$ respectively, and $\sigma_{i}$ are the nonnegative singular values in nonincreasing order. The leading singular triplet is known explicitly: $\sigma_{1}=1$, $u_{1}=\mathbbm{1}_{\mathcal{Z}}$ and $v_{1}=\mathbbm{1}_{\mathcal{X}}$ corresponding to the constant functions. If $\mathcal{T}$ is infinite-rank, then $\sigma_{i}\to 0$ as $i\to\infty$. Under the assumption that $\mathcal{T}$ is injective (Assumption˜1), we have $\mathcal{N}(\mathcal{T})=\{0\}$, hence $\{v_{i}\}_{i\geq 1}$ forms an orthonormal basis (ONB) of $L_{2}(X)$ and $\sigma_{i}>0$ for all $i\geq 1$. Finally, since $\mathcal{T}$ is Hilbert–Schmidt, $\sum_{i\geq 1}\sigma_{i}^{2}<\infty$. For any $d\geq 1$, define the leading eigensubspaces of $\mathcal{T}$ as
@@
In addition, define $\mathcal{T}_{d}\doteq\sum_{i=1}^{d}\sigma_{i}u_{i}\otimes v_{i}$ the rank-$d$ truncation of $\mathcal{T}$. To avoid ambiguity in the definition of $\mathcal{U}_{d},\mathcal{V}_{d}$ and $\mathcal{T}_{d}$, we assume that $\sigma_{d}>\sigma_{d+1}$.
@@ 3.1 Sieve 2SLS estimator
possibly subject to regularization. In the algorithm 2SLS, we consider features on both $\mathcal{X}$ and $\mathcal{Z}$, and both steps (estimating $h\mapsto\mathcal{T}h$ and solving Eq.˜4) are carried out via linear regression in feature space. The two sets of features may be infinite-dimensional—e.g., using Reproducing Kernel Hilbert Spaces (RKHS)—or finite-dimensional—e.g., using splines, wavelets, or neural network features. When the feature dimension grows with sample size, we obtain a sieve. Since our focus is on neural network-based features, we present the finite-dimensional version, noting that the ideas naturally extend to infinite-dimensional cases with proper regularization [33, 29]. We refer to Appendix B for a description of other 2SLS methods.
@@
3.1.0.0.1 2SLS versus saddle-point estimator.
@@ 3.2 Convergence rates for Sieve 2SLS
The first term captures the approximation error, while the second term corresponds to the estimation error arising from finite samples. When standard bases such as cosine, spline, or wavelet functions are used, the approximation error is well-understood [3, 6]. If the sieve ill-posedness $\tau_{\varphi,d}$ grows at a polynomial or exponential rate in $d$, one can choose $d=d(n)$ to balance the two terms optimally. Under such growth conditions and smoothness assumptions on $h_{0}$ (e.g., $h_{0}$ belongs to a Sobolev ball), the resulting convergence rate is minimax optimal.
@@ 3.3 Convergence rates for Sieve 2SLS with spectral features
Corollary 1 (Sieve 2SLS with spectral features).
@@
We postpone the proof to Appendix A.2.
The good, the bad, and the ugly. The bound in Corollary 1 decomposes the generalization error into two terms: spectral alignment, encoded in the approximation error $\left\|(\Pi_{\mathcal{X},d})_{\perp}h_{0}\right\|$, which measures how well $h_{0}$ lies in the top-$d$ singular space; and ill-posedness, captured by the inverse singular value $\sigma_{d}^{-1}$. In the good regime, $h_{0}$ lies mostly in $\mathcal{U}_{d}$ and $\sigma_{d}$ decays slowly, so both terms are small. In the bad regime, alignment is favorable but fast spectral decay inflates the estimation error. In the ugly regime, $h_{0}$ is misaligned with the top eigenspaces, so the approximation error dominates and estimation fails regardless of $\sigma_{d}$ or $n$.
@@ 4 Sieve 2SLS with learned spectral features
where the last term is independent of $(\varphi,\psi)$. We provide a proof of this equivalence in Appendix A.3. The population loss can be estimated from samples $\tilde{\mathcal{D}}_{m}=\{(\tilde{x}_{i},\tilde{z}_{i})\}_{i=1}^{m}$ as
@@
The spectral contrastive loss is the basis of the recent algorithm proposed in [34], where $\varphi$ and $\psi$ are parametrized by neural networks and optimized via stochastic gradient descent. While [34] motivate this objective heuristically, Eq.˜6 shows that it directly targets the leading singular structure of $\mathcal{T}$. The use of objectives similar to Eq.˜6 for operator learning or Eq.˜7 for representation learning has a rich history that we detail in Appendix B.
@@
We argue that this assumption is overly strong: it holds if and only if $\mathcal{T}$ has rank at most $d$, in which case the spectral features perfectly capture the structure of $\mathcal{T}$. This result is formalized in ˜5, Appendix A.3. We should instead interpret Eq. (9) as an approximation rather than a literal assumption. In practice, the learned features $\varphi,\psi$ provide a rank-$d$ approximation of $\mathcal{T}$, and the quality of this approximation governs the performance of the resulting Sieve 2SLS estimator.
@@
Theorem 3.
Let Assumptions˜1, 2, 4 and 5 hold and let $\hat{\varepsilon}_{d}\doteq\left\|\hat{\mathcal{T}}_{d}-\mathcal{T}_{d}\right\|_{\mathrm{op}}$ be such that $\hat{\varepsilon}_{d}<(1-1/\sqrt{2})\sigma_{d}$.
@@
We postpone the proof to Appendix A.3. The result quantifies how close the learned features must be to $\mathcal{T}_{d}$ in order to preserve the desirable properties of spectral 2SLS. Theorem˜3-i) shows that as long as the learned operator $\hat{\mathcal{T}}_{d}$ is a good approximation of the true rank-$d$ truncation $\mathcal{T}_{d}$, the learned features inherit a favorable measure of ill-posedness. Theorem˜3-ii) further ensures that if contrastive learning yields a near-optimal rank-$d$ approximation of $\mathcal{T}$, the downstream Sieve 2SLS estimator retains strong statistical guarantees. Importantly, the spectral features generalization error $\hat{\varepsilon}_{d}$ encapsulates the statistical and optimization errors in learning spectral features. It depends on the number of samples used to train the contrastive loss, the expressivity and architecture of the neural networks parametrizing $\varphi$ and $\psi$, and the complexity of the leading eigensubspaces of $\mathcal{T}$. As $\mathcal{L}_{d}(\varphi,\psi)$ converges to $\mathcal{L}_{d}^{\mathrm{min}}\doteq\left\|\mathcal{T}-\mathcal{T}_{d}\right\|_{\scriptscriptstyle\mathrm{HS}}^{2},$ $\hat{\varepsilon}_{d}$ converges to 0 by Theorem˜2. Precisely controlling $\hat{\varepsilon}_{d}$ remains an important question for future work.
Figure 1: Dependence of the MSE of instrumental variable regression on the decay rates of the spectrum and coefficients of the structural function. IV fitting and the MSE evaluations were repeated 500 times per parameter set, rendering the standard error of these estimates negligible.
@@ 5 Experiments
Let $\pi_{XZ}$ be a probability distribution on $\mathcal{X}\times\mathcal{Z}$ with marginals $\pi_{X}$ and $\pi_{Z}$. Let $\mathcal{T}:L_{2}(\mathcal{X},\pi_{X})\to L_{2}(\mathcal{Z},\pi_{Z})$ be the conditional expectation operator associated with $\pi_{XZ}$ and let $h_{0}\in L_{2}(\mathcal{X},\pi_{X})$. Sample $(X,Z)\sim\pi_{XZ}$, $V\sim\mathcal{N}(0,\sigma^{2})$ and set $U=\mathcal{T}h_{0}(Z)-h_{0}(X)+V$ and $Y=h_{0}(X)+U$. Then $(U,Z,X,Y)$ is a sample from the NPIV model of Eq. (1).
Let $d=11$, $\mathcal{X}=[0,2\pi]$, $\mathcal{Z}=[0,2\pi]$, $\sigma^{2}=0.1$, $f(x)=(\sin(x),\sin(2x),\ldots,\sin((d-1)x))$, $g(z)=(\sin(z),\sin(2z),\ldots,\sin((d-1)z))$, and let $P,Q$ be two orthogonal $(d-1)\times(d-1)$ matrices. Define $v:\mathcal{X}\to\mathbb{R}^{d-1}$ and $u:\mathcal{Z}\to\mathbb{R}^{d-1}$ as $v(x)=Pf(x)$ and $u(z)=Qg(z)$. We set $\mathcal{T}=\mathbbm{1}_{\mathcal{Z}}\otimes\mathbbm{1}_{\mathcal{X}}+\sum_{i=1}^{d-1}\sigma_{i}u_{i}\otimes v_{i}$. $(X,Z)$ is jointly sampled with rejection sampling. $Z,X$ are uniform on $0,2\pi]$ and $X$ given $Z$ admits $\mathcal{T}$ as conditional expectation operator. We then sample $Y$ following ˜3. The multiplication of the trigonometric functions by the orthogonal matrices $P,Q$ ensures that the consecutive singular functions of $\mathcal{T}$ are comparably complex. By keeping the difficulty constant per singular function we can isolate the influence of the singular value decay on the performance of the final estimator. We set the singular values $\{\sigma_{i}\}_{i=1}^{d-1}$ to decay linearly, from some initial value $\sigma_{1}$ to a $\sigma_{d-1}=c_{\sigma}\sigma_{1}$, $c_{\sigma}\in[0,1]$. To control spectral alignment, we set $h_{0}=\sum_{i=1}^{d-1}\alpha_{i}v_{i}$ subject to $\|\alpha\|_{\ell_{2}}=1$, and allow $\{\alpha_{i}\}_{i=1}^{d-1}$ to decay linearly from $\alpha_{1}$ to $\alpha_{d-1}=c_{\alpha}\alpha_{1}$, $c_{\alpha}\in[0,1]$. Each value of $c_{\alpha}\in[0,1]$ controls the alignment of $h_{0}$ with the top singular functions of $\mathcal{T}$: larger values correspond to stronger alignment. Similarly, $c_{\sigma}\in[0,1]$ controls the ill-posedness of the problem: smaller values indicate faster decay of singular values and hence more severe ill-posedness.
@@
In Figure˜1, we report the mean squared error (MSE) of two 2SLS estimators across a range of spectral decay rates ($c_{\sigma}$) and structural function decay rates ($c_{\alpha}$). Left: Oracle spectral features: The left panel uses 2SLS with fixed features $f$ and $g$ that form spectral features. This represents an oracle setting where feature learning is not needed. As expected, the MSE decreases from top to bottom as the problem becomes better conditioned (i.e., as $c_{\sigma}$ increases). Additionally, the MSE decreases from left to right, as smaller values of $c_{\alpha}$ correspond to stronger spectral alignment: more of $h_{0}$ is concentrated on the top singular functions of $\mathcal{T}$. These trends are fully consistent with the bound established in Corollary˜1. Right: Learned features. In the right panel, we learn 50 features from data by minimizing the empirical contrastive loss (8) using a neural network. We then build a 2SLS estimator on top of the learned features. The same qualitative trends are observed: MSE decreases with slower singular value decay and better alignment. Moreover, when $c_{\sigma}$ is small (top rows), learning the entire eigenspaces becomes difficult. In this regime, performance is most sensitive to $c_{\alpha}$. In particular, if $h_{0}$ is concentrated on the top singular functions (small $c_{\alpha}$), good recovery is still possible. In contrast, when $c_{\sigma}$ is large (bottom rows), the singular functions are equally important and more easily learned, rendering the alignment of $h_{0}$ less critical. The mismatch between the oracle and learned settings is because we do not exactly recover the singular feature spaces; as predicted by Theorem˜3. It is also worth noting that when $h_{0}$ is concentrated on high singular values, it may sometimes be beneficial for the singular values to decay quickly, by making it easier for the feature-learning model to learn the top singular functions well. This may explain the slight top-bottom increase in losses on the right end of the right panel.

### v3

@@ Abstract
We address the problem of causal effect estimation in the presence of hidden confounders, using nonparametric instrumental variable (IV) regression. A leading strategy employs spectral features - that is, learned features spanning the top eigensubspaces of the operator linking treatments to instruments. We derive a generalization error bound for a two-stage least squares estimator based on spectral features, and gain insights into the method’s performance and failure modes. We show that performance depends on two key factors, leading to a clear taxonomy of outcomes. In a good scenario, the approach is optimal. This occurs with strong spectral alignment, meaning the structural function is well-represented by the top eigenfunctions of the conditional operator, coupled with this operator’s slow eigenvalue decay, indicating a strong instrument. Performance degrades in a bad scenario: spectral alignment remains strong, but rapid eigenvalue decay (indicating a weaker instrument) demands significantly more samples for effective feature learning. Finally, in the ugly scenario, weak spectral alignment causes the method to fail, regardless of the eigenvalues’ characteristics. Our synthetic experiments empirically validate this taxonomy. We further introduce a practical procedure to estimate these spectral properties from data, allowing practitioners to diagnose which regime a given problem falls into. We apply this method to the dSprites dataset, demonstrating its utility.
@@ 1 Introduction
$\bullet$ We empirically validate the “good–bad–ugly” taxonomy through synthetic experiments designed to isolate the effects of spectral alignment and singular value decay on generalization error. We further introduce a practical, data-driven procedure to estimate these key spectral properties (spectral alignment and singular value decay), allowing practitioners to diagnose their problem’s regime. We demonstrate this procedure on the dSprites dataset [30].
@@
Features can be learned in various ways. One line of work learns them jointly with the regression objective, as in [16, 23, 33, 40], who propose end-to-end frameworks that simultaneously optimize over the feature representations and second-stage parameters. In contrast, [36, 38] propose to learn features that reflect the eigenstructure of $\mathcal{T}$, independently of the outcome $Y$, using a spectral contrastive loss. This may appear suboptimal, since incorporating $Y$ could, in principle, yield more informative representations, however it avoids the optimization difficulties inherent to nonconvex, joint objectives involving all three variables $(X,Y,Z)$. This decoupled spectral approach was shown in certain experimental settings to outperform end-to-end alternatives, as demonstrated in [36]. Our work places this empirical finding in context, since it likely arises due to the experiments in that work satisfying the “good” regime in the present work. We provide further discussion in Appendix D.
Structure of the paper. Section 2 introduces the notation and preliminaries. Section 3 reviews the Sieve 2SLS estimator and presents generalization bounds highlighting the role of spectral features. Section 4 shows how to learn such features via a contrastive loss and connects this to the method of [36]. Section˜5 presents a practical procedure to estimate the spectral properties from data, enabling an empirical diagnosis of our taxonomy. Section 6 provides synthetic experiments, including an application to the dSprites dataset, that validate our theoretical taxonomy. Proofs are deferred to the Appendix.
@@ 2.2 On the Conditional Expectation Operator: NPIV as an Ill-posed Inverse Problem
The existence part is essential: without it, there is no structural function $h_{0}$ satisfying the model in Eq.˜1. When $\mathcal{T}$ is not injective, we lose identifiability (i.e., unicity of the solution). In such scenarios, it is possible to establish consistency to a minimum-norm solution of the inverse problem Eq.˜2, namely $h_{*}\doteq\mathcal{T}^{\dagger}r_{0}$ [15]. [^2] This corresponds to the unique element in the pre-image of $r_{0}$ under $\mathcal{T}$ that minimizes the $L_{2}$-norm and is equivalently characterized as the unique element of the set $\mathcal{T}^{-1}(\{r_{0}\})\cap\mathcal{N}(\mathcal{T})^{\perp}$ [2]. We assume injectivity henceforth to simplify the discussion.
@@
{restatable}
propTHS Under Assumption˜2, $\mathcal{T}$ is a Hilbert–Schmidt operator and thus compact. This result is well known in the inverse problems literature; we provide a proof in Appendix C.1.
@@
where $\{u_{i}\}$ and $\{v_{i}\}$ are orthonormal basis of $L_{2}(Z)$ and $L_{2}(X)$ respectively (left and right singular functions, respectively), and $\sigma_{i}$ are the nonnegative singular values in nonincreasing order. The leading singular triplet is known explicitly: $\sigma_{1}=1$, $u_{1}=\mathbbm{1}_{\mathcal{Z}}$ and $v_{1}=\mathbbm{1}_{\mathcal{X}}$ corresponding to the constant functions. Under the assumption that $\mathcal{T}$ is injective (Assumption˜1) $\sigma_{i}>0$ for all $i\geq 1$. Finally, since $\mathcal{T}$ is Hilbert–Schmidt, $\sum_{i\geq 1}\sigma_{i}^{2}<\infty$. For any $d\geq 1$, define the leading eigensubspaces of $\mathcal{T}$ as
@@
In addition, define $\mathcal{T}_{d}\doteq\sum_{i=1}^{d}\sigma_{i}u_{i}\otimes v_{i}$ the rank-$d$ truncation of $\mathcal{T}$. To avoid ambiguity in the definition of $\mathcal{U}_{d},\mathcal{V}_{d}$ and $\mathcal{T}_{d}$, we assume that $\sigma_{d}>\sigma_{d+1}$. We refer to Appendix˜A for more details on spectral theory.
@@ 3.1 Sieve 2SLS estimator
possibly subject to regularization. In the algorithm 2SLS, we consider features on both $\mathcal{X}$ and $\mathcal{Z}$, and both steps (estimating $h\mapsto\mathcal{T}h$ and solving Eq.˜4) are carried out via linear regression in feature space. The two sets of features may be infinite-dimensional—e.g., using Reproducing Kernel Hilbert Spaces (RKHS)—or finite-dimensional—e.g., using splines, wavelets, or neural network features. When the feature dimension grows with sample size, we obtain a sieve. Since our focus is on neural network-based features, we present the finite-dimensional version, noting that the ideas naturally extend to infinite-dimensional cases with proper regularization [35, 31]. We refer to Appendix D for a description of other 2SLS methods.
@@
2SLS versus saddle-point estimator.
@@ 3.2 Convergence rates for Sieve 2SLS
The first term captures the approximation error, while the second term corresponds to the estimation error arising from finite samples. When standard bases such as cosine, spline, or wavelet functions are used, the approximation error is well-understood. If the sieve ill-posedness $\tau_{\varphi,d}$ grows at a polynomial or exponential rate in $d$, one can choose $d=d(n)$ to balance the two terms optimally [3]. Under such growth conditions and smoothness assumptions on $h_{0}$ (e.g., $h_{0}$ belongs to a Sobolev ball), the resulting convergence rate is minimax optimal [8].
@@ 3.3 Convergence rates for Sieve 2SLS with spectral features
{restatable}
[Sieve 2SLS with spectral features]corollaryratespectral
@@
We postpone the proof to Appendix C.2.
The good, the bad, and the ugly. The bound in Definition˜2 decomposes the generalization error into two terms: spectral alignment, encoded in the approximation error $\left\|(\Pi_{\mathcal{X},d})_{\perp}h_{0}\right\|$, which measures how well $h_{0}$ lies in the top-$d$ singular space; and ill-posedness, captured by the inverse singular value $\sigma_{d}^{-1}$. In the good regime, $h_{0}$ lies mostly in $\mathcal{U}_{d}$ and $\sigma_{d}$ decays slowly, so both terms are small. In the bad regime, alignment is favorable but fast spectral decay inflates the estimation error. In the ugly regime, $h_{0}$ is misaligned with the top eigenspaces, so the approximation error dominates and estimation fails regardless of $\sigma_{d}$ or $n$. In Appendix˜B, we show that the bound in Corollary 1 is tight under a source condition and a singular value decay condition.
@@ 4 Sieve 2SLS with learned spectral features
where the last term is independent of $(\varphi,\psi)$. We provide a proof of this equivalence in Appendix C.3. The population loss can be estimated from samples $\tilde{\mathcal{D}}_{m}=\{(\tilde{x}_{i},\tilde{z}_{i})\}_{i=1}^{m}$ as
@@
The spectral contrastive loss is the basis of the recent algorithm proposed in [36], where $\varphi$ and $\psi$ are parametrized by neural networks and optimized via stochastic gradient descent. While [36] motivate this objective heuristically, Eq.˜6 shows that it directly targets the leading singular structure of $\mathcal{T}$. The use of objectives similar to Eq.˜6 for operator learning or Eq.˜7 for representation learning has a rich history that we detail in Appendix D.
@@
We argue that this assumption is overly strong: it holds if and only if $\mathcal{T}$ has rank at most $d$, in which case the spectral features perfectly capture the structure of $\mathcal{T}$. This result is formalized in ˜4, Appendix C.3. We should instead interpret Eq. (9) as an approximation rather than a literal assumption. In practice, the learned features $\varphi,\psi$ provide a rank-$d$ approximation of $\mathcal{T}$, and the quality of this approximation governs the performance of the resulting Sieve 2SLS estimator.
@@
{restatable}
theoremmeasillpossievencp Let Assumptions˜1, 2, 4 and 5 hold and let $\hat{\varepsilon}_{d}\doteq\left\|\hat{\mathcal{T}}_{d}-\mathcal{T}_{d}\right\|_{\mathrm{op}}$ be such that $\hat{\varepsilon}_{d}<(1-1/\sqrt{2})\sigma_{d}$.
@@
We postpone the proof to Appendix C.3. The result quantifies how close the learned features must be to $\mathcal{T}_{d}$ in order to preserve the desirable properties of spectral 2SLS. Assumption˜5-i) shows that as long as the learned operator $\hat{\mathcal{T}}_{d}$ is a good approximation of the true rank-$d$ truncation $\mathcal{T}_{d}$, the learned features inherit a favorable measure of ill-posedness. Assumption˜5-ii) further ensures that if contrastive learning yields a near-optimal rank-$d$ approximation of $\mathcal{T}$, the downstream Sieve 2SLS estimator retains strong statistical guarantees. Importantly, the spectral features generalization error $\hat{\varepsilon}_{d}$ encapsulates the statistical and optimization errors in learning spectral features. It depends on the number of samples used to train the contrastive loss, the expressivity and architecture of the neural networks parametrizing $\varphi$ and $\psi$, and the complexity of the leading eigensubspaces of $\mathcal{T}$. As $\mathcal{L}_{d}(\varphi,\psi)$ converges to $\mathcal{L}_{d}^{\mathrm{min}}\doteq\left\|\mathcal{T}-\mathcal{T}_{d}\right\|_{\scriptscriptstyle\mathrm{HS}}^{2},$ $\hat{\varepsilon}_{d}$ converges to 0 by Theorem˜2. Precisely controlling $\hat{\varepsilon}_{d}$ lies beyond the scope of this paper and remains an important question for future work.
5 Estimating alignment and spectral decay from data
Having established the good-bad-ugly taxonomy based on spectral properties and its interaction with the structural function $h_{0}$, we now discuss how to estimate these spectral properties from data. Namely, we present a methodology to estimate the spectral decay of the operator $\mathcal{T}$ as well as spectral alignment with the structural function $h_{0}$. Recall that $\mathcal{T}$ is compact and admits an SVD of the form Eq.˜3, with the key relationship that $\sigma_{i}v_{i}=\mathcal{T}^{*}u_{i}$. Therefore, for all $i\geq 1$, we have
$\langle v_{i},h_{0}\rangle_{L_{2}(X)}=\frac{1}{\sigma_{i}}\langle\mathcal{T}^{*}u_{i},h_{0}\rangle_{L_{2}(X)}=\frac{1}{\sigma_{i}}\langle u_{i},r_{0}\rangle_{L_{2}(Z)}=\frac{1}{\sigma_{i}}\mathbb{E}[Y\cdot u_{i}(Z)],$ (10)
where we used $\mathcal{T}h_{0}=r_{0}$, the definition of $r_{0}$ and the tower property of the conditional expectation. This relationship shows that the alignment coefficients $\langle v_{i},h_{0}\rangle_{L_{2}(X)}$ can be estimated from data if we have estimates of the singular functions $u_{i}$ and singular values $\sigma_{i}$. In Section˜5, we provide a practical estimator for Eq.˜10 using learned features and derive its estimation guarantees, which depend on the operator error $\hat{\varepsilon}_{d}$.
We apply this procedure in Section˜6 to diagnose the spectral properties of the dSprites dataset [30], a popular benchmark in the nonparametric IV and proxy literature [40, 41, 36].
Figure 1: Dependence of the MSE of instrumental variable regression on the decay rates of the spectrum and coefficients of the structural function. IV fitting and the MSE evaluations were repeated 500 times per parameter set, rendering the standard error of these estimates negligible. Left: Oracle spectral features; Right: Learned features.
@@ 6.1 Synthetic datasets
Let $\pi_{XZ}$ be a probability distribution on $\mathcal{X}\times\mathcal{Z}$ with marginals $\pi_{X}$ and $\pi_{Z}$. Let $\mathcal{T}:L_{2}(\mathcal{X},\pi_{X})\to L_{2}(\mathcal{Z},\pi_{Z})$ be the conditional expectation operator associated with $\pi_{XZ}$ and let $h_{0}\in L_{2}(\mathcal{X},\pi_{X})$. Sample $(X,Z)\sim\pi_{XZ}$, $V\sim\mathcal{N}(0,\sigma^{2})$ and set $U=\mathcal{T}h_{0}(Z)-h_{0}(X)+V$ and $Y=h_{0}(X)+U$. Then $(U,Z,X,Y)$ is a sample from the NPIV model of Eq.˜1.
Let $d=11$, $\mathcal{X}=[0,2\pi]$, $\mathcal{Z}=[0,2\pi]$, $\sigma^{2}=0.1$, $f(x)=(\sin(x),\sin(2x),\ldots,\sin((d-1)x))$, $g(z)=(\sin(z),\sin(2z),\ldots,\sin((d-1)z))$, and let $P,Q$ be two orthogonal $(d-1)\times(d-1)$ matrices. Define $v:\mathcal{X}\to\mathbb{R}^{d-1}$ and $u:\mathcal{Z}\to\mathbb{R}^{d-1}$ as $v(x)=Pf(x)$ and $u(z)=Qg(z)$. We set $\mathcal{T}=\mathbbm{1}_{\mathcal{Z}}\otimes\mathbbm{1}_{\mathcal{X}}+\sum_{i=1}^{d-1}\sigma_{i}u_{i}\otimes v_{i}$. $(X,Z)$ is jointly sampled with rejection sampling. $Z,X$ are uniform on $0,2\pi]$ and $X$ given $Z$ admits $\mathcal{T}$ as conditional expectation operator. We then sample $Y$ following ˜2. The multiplication of the trigonometric functions by the orthogonal matrices $P,Q$ ensures that the consecutive singular functions of $\mathcal{T}$ are comparably complex. By keeping the difficulty constant per singular function we can isolate the influence of the singular value decay on the performance of the final estimator. We set the singular values $\{\sigma_{i}\}_{i=1}^{d-1}$ to decay linearly, from some initial value $\sigma_{1}$ to a $\sigma_{d-1}=c_{\sigma}\sigma_{1}$, $c_{\sigma}\in[0,1]$. To control spectral alignment, we set $h_{0}=\sum_{i=1}^{d-1}\alpha_{i}v_{i}$ subject to $\|\alpha\|_{\ell_{2}}=1$, and allow $\{\alpha_{i}\}_{i=1}^{d-1}$ to decay linearly from $\alpha_{1}$ to $\alpha_{d-1}=c_{\alpha}\alpha_{1}$, $c_{\alpha}\in[0,1]$. Each value of $c_{\alpha}\in[0,1]$ controls the alignment of $h_{0}$ with the top singular functions of $\mathcal{T}$: larger values correspond to weaker alignment. Similarly, $c_{\sigma}\in[0,1]$ controls the ill-posedness of the problem: smaller values indicate faster decay of singular values and hence more severe ill-posedness.
@@
In Figure˜1, we report the mean squared error (MSE) of two 2SLS estimators across a range of spectral decay rates ($c_{\sigma}$) and structural function decay rates ($c_{\alpha}$). The left panel uses 2SLS with fixed features $f$ and $g$ that form the exact spectral features. This represents an oracle setting. The MSE decreases from top to bottom as the problem becomes better conditioned (i.e., as $c_{\sigma}$ increases). Additionally, the MSE decreases from left to right, as smaller values of $c_{\alpha}$ correspond to stronger spectral alignment: more of $h_{0}$ is concentrated on the top singular functions of $\mathcal{T}$. These trends are fully consistent with the bound established in Definition˜2. In the right panel, we learn 50 features from data by minimizing the empirical contrastive loss Eq.˜8 using a neural network. We then build a 2SLS estimator on top of the learned features. The same qualitative trends are observed: MSE decreases with slower singular value decay and better alignment. Moreover, when $c_{\sigma}$ is small (top rows), learning the entire eigenspaces becomes difficult. In this regime, performance is most sensitive to $c_{\alpha}$. In particular, if $h_{0}$ is concentrated on the top singular functions (small $c_{\alpha}$), good recovery is still possible. In contrast, when $c_{\sigma}$ is large (bottom rows), the singular functions are equally important and more easily learned, rendering the alignment of $h_{0}$ less critical. The mismatch between the oracle and learned settings is because we do not exactly recover the singular feature spaces; as predicted by Assumption˜5. It is also worth noting that when $h_{0}$ is concentrated on high singular values, it may sometimes be beneficial for the singular values to decay quickly, by making it easier for the feature-learning model to learn the top singular functions well. This may explain the slight top-bottom increase in losses on the right end of the right panel.